OpenStudy (anonymous):

How long will it take for $8000 to grow to $45,300 at an interest rate of 11% if the interest is compounded continuously? Round the number of years to the nearest hundredth

OpenStudy (anonymous):

I am presuming here that it is compounded annually (as nothing is mentioned about it) A = P (1 + R)ⁿ HERE A = 45,300 P = 8000 R = 11% = 0.11 n1 = ????? 45,300 = 8000 (1 + 0.11)ⁿ 45,300/8000 = (1.11)ⁿ 5.6625 = (1.11)ⁿ Now 1.11^16.614 = 5.66234034 so n = 16.614 or 16.61 years

OpenStudy (anonymous):

For me the question was not very clear as nothing was mentioned about the compounding period -yearly, half-yearly or quarterly.. I hv calculated assuming that it is yearly.....